A Bibliography of Serif: The Magazine of Type and Typography

Nelson H. F. Beebe Department of Mathematics

University of Utah Salt Lake City, UT 84112

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)

WWW URL: http://www.math.utah.edu/~beebe/

12 August 2014 Version 1.08

Title word cross-reference

’94 [Ano95d]. ’95 [Ano95e, Ano96b, Ano95e].

a-changing [Pri95]. Adobe [BM98, Hos95a, Hos95b]. Adrian [Hun98]. age [Bri95b]. Alan [Hos95d]. Aldo [Ano96b]. Alexander [Pas94]. alphabet [Hos95h]. Alphabets [BM98]. American [Ano94b, Hos95c, Sax94]. Anatomy [Pas94]. Apologies [Pri95]. Aries [Ano96d]. Arrangement [Hos96a]. Arrighi [Hos94b]. Art [Hos96a]. ATF [Ano95d]. Atlas [Hos95d]. ATypI [Ano95d, Ano95e, BM98, Ano96b, BM98]. award [Ano95e].

baroque [Bri95b]. Bartram [Hos95d]. Baskerville [Bri96]. Baudin [Ric94]. be [Bar95]. Beata [Ano98]. beginning [Bea95].

Bell [Bri96]. Big [Ano94a]. bitmap [Ano96b]. Bitstream [Ano95e]. Book [Ano94b, Ano95a, Ano95b, Ano95c, Hos96a, Hos94a, Pas94, Ric94, Hos96d]. Books [Hos95e, Hos95c, Hos95b, Hos95d, Sax94, Hos95b]. Borek [Ano96d]. Briefly [Ano94b, Ano95d, Ano95e, Ano96b, BM98]. Bulmer [Ano95g]. Bureau [Ano95e]. Butterick [Hos96d].

Callifonts [Bar95]. Can [Bar95]. Carl [Ano95e]. Carlos [Ano96c]. Carol [Hos95f]. Carter [Lan94a]. Cash [Hal98]. Caslon [Ano94a, Hos95i]. Casual [Ano94a]. Catalog [Ano94b]. Catalogs [Ano95e]. Catalogue [Hos95e]. CD [Ano96b, Hos96c]. CDs [Ano94b]. Centaur [Hos94b]. Center [Ano95d]. centuries [Joh95]. Century [Ano95c, Reh96, Big94, Hos95c]. challenges [Bea96]. Change [Kee98]. changing [Pri95, Sch94]. Chelsea [Ano94a]. Circle [Ano96b]. Civilité [Ano96d]. Clarety [Swa94]. classification [Bri94]. Cleary [Ano94b]. collage [Ano96c]. Comments [WR98]. Company [Ano96b]. Competition [Ano96a]. Cone [Lan94a]. Conference [BM98]. Contemporary [Hos95e, Hos95c, Hos95b, Hos95d]. Conversation [Ano96c, Hos95f]. Cool [Hal98]. Copal [Ano95f]. creating [Reh96]. Credit [Ano95b]. Cresci [Ano98]. Cruz [Ano94b]. crystal [Swa94].

David [Lon95]. Dealing [Bea96]. death [Ano95d]. debate [Ano96b]. Design [Hos96a, Ano96a, Fra98, Lan95, Bea96, Car94, Lan94b]. designed [Hos96d]. designers [Bri96, Ano95c]. Designing [Bea96]. development [Bri96]. Digital [Lan94a]. discuss [Ano96c]. Donatello [Ano98]. Drinking [Swa94]. Duckling [Ano96d].

Ed [Ano94b]. Education [Fra98]. Elroy [Ano94a]. Emigre [Ano94b]. Endpapers [Hos94c, Hos95g, Hos96b, Hos98]. enough [Lan95]. entering [Ano96a]. Essays [Hos96a]. examination [Bri96]. expansion [Ano96b]. exposition [Ano96b].

face [Hos96d]. Faces [Ano98, Ano95f, Ano95g, Ano96d]. fashion [Lan95]. Fellowship [Ano94b]. Fernand [Ric94]. few [Lub95]. FF [Ano94a, Ano95f]. FH [Ano96d]. fifteenth [Hos95h]. fine [Sol94]. Finer [Hos96a]. first [Reh96]. Fleischmann [Ano95a]. floppy [Ano95d]. Font [BM98, Ano95e, Bar95, Ano94b]. FontHaus [Ano96b]. Fonts [Hal98]. Form [Hos96a]. foundry [Lan94a, Sch96]. Fournier [Bri96]. France [BM98]. Frutiger [Hun98]. FUSE [Ano95e].

Galápagos [Lan95]. Ghiberti [Ano98]. goblet [Swa94]. Goes [Sch96]. Good [Hos96a, Bar95]. Goudy [Ano95d]. Group [Lan95]. grunge [Ano96c, Sch96]. guard [Sch94]. Guide [Reh96]. GX [Ano94b].

Hairspray [Ano95g]. Hand

[Bri96, Bri95a, Sol94]. Handsome [Ano95g]. Headliners [Ano94b]. height [Ano96b]. Herzog [Ano95e]. Hitchhiker [Reh96]. Holding [Bri95a]. Horizontalité [Por98]. Hot [Hal98]. House [Sch96]. Human [Fra98].

ideas [Bri95a]. Inc. [Lan94a]. Industries [Sch96]. influences [Car94]. Information [Ano96a, Ano95d, Ano95e]. Initials [Ano95g]. ink [Bri95b]. Instanter [Ano95f]. Interaction [Fra98]. International [Ano94b, Ano95d, Hos96c]. Introduction [Hos96a]. Invisible [Bri96]. italic [Bri95b, Sol94]. ITC [Ano96b].

Jamboree [BM98]. James [Hos95d]. Jecho [Ano96d].

Kalendar [Ano96d]. Kindersley [Lon95]. Kinross [Hos94a].

Lawson [Pas94]. lawsuit [Ano95e].

legibility [Lub95]. Letter [Bea96, Hos95h]. Letterforms [Bri96, Bri94, Bri95a]. Lettering [Fra98]. Letterpress [Reh96, Due95]. Letters [Ano95b, WR98, Lon95]. look [Due95, Hos96d, Sch96]. love [Lon95]. Lucida [Ano94a]. Lyon [BM98].

M&H [Hos95e]. Mac [Hos95c]. made [Bar95]. Magneto [Ano95f]. Massachusetts [BM98]. Masters [Ano94b]. Matthew [Hos96d]. McGrew [Hos95c]. Metal [Due95, Hos95c]. metaphysics [Bri95a]. methods [Reh96]. Microsoft [BM98]. MJ [Ano96d]. Modern [Hos94a].

Monotype [Ano95g, Ano96b]. Morality [Hos96a]. Morris [Pet94]. Mr. [Pri95]. much [Lan95]. Multiple [Ano94b].

Neighborhood [Sch96]. Neoclassical [Bri96]. Neonstream [Ano95f]. News [Ano95d]. Noted [Ano94b, Ano95d, Ano95e, Ano96b, BM98]. Novarese [Ano96b]. November [BM98].

O [Hos95h]. October [BM98]. One [Bri96, Ano96d]. OpenType [BM98]. originality [Ano96c]. overview [Reh96]. own [Sch96].

Painting [Bri95b]. Paper [Bea96]. Part [Bri96]. party [Ano95e]. physics [Bri95a]. Pietra [Ano98]. Poetica [Hos95a]. Points [Hos96a]. Pontif [Ano98]. Precision [Ano95e]. printing [Due95, Reh96]. producing [Sch96]. publications [Ano95d]. punchcutting [Car94, Ano95a].

Quickdraw [Ano94b].

Raceway [Ano95f]. Readability [WR98, Ano96c, Lub95]. Reader [WR98]. Reading [BM98]. redux [Ano95e]. renaissance [Bri95a]. Restructures [BM98]. return [Ano95d, Ano96b]. reviews [Ano95a, Ano95b, Ano95c, Hos96a, Hos94a, Pas94, Ric94]. revival [Ano95e]. Robin [Hos94a]. Rocket [Ano95f]. roman [Bri95b].

Santa [Ano94b]. Schneidler [Ano95g]. Schoolbook [Big94]. Segura [Ano96c]. serif [Car94]. sheet [Bea96]. Show [Ano94b]. Simona [Ano98]. Spacing [Hos96a]. Specimen [Hos95e, Hos95c, Hos95b, Hos95d, Sax94]. specimens [Sax95]. state [Due95]. steel [Bri95b]. style [Lub95, Sch96]. Subiaco [Ano95e]. such [Bri96]. superhighway [Ano95d, Ano95e]. Supermarket [Van98].

Sutton [Hos95d]. symposium [Ano96b].

T [Lan94b]. T-26 [Lan94b]. Technology [Ano95h, Hos96c, Joh95]. Techo [Ano96d]. Theory [Van98]. There [Sch96]. Thoughts [Big94, Lub95]. times [Pri95]. Too [Lan95]. Toronto [Ano95e]. Tracy [Ano95e]. tradition [Lub95]. Treacyfaces [Ano94b]. Twentieth [Ano95c, Hos95c]. Twenty [Reh96]. Twenty-first [Reh96]. Twist [Ano94a]. Two [Joh95, Bar95]. Twombly [Hos95f]. Type [Ano95d, Ano95e, Ano95c, Hos96a, Hos95e, Hos95c, Hos95b, Hos95d, Kee98, Car94, Due95, Joh95, Lan94a, Reh96, Sax94, Sax95]. Typecasting [Ano94b]. TypeCon [BM98]. Typeface [Pas94]. Typefaces [Hos95c]. Typeforms [Hos95d]. Typefounders [Hos96c]. types [Bri95b, Pet94]. Typo.icon.ography [Hos96a]. Typographic [Ano96b, Ano96a, Hun98, Ano96c, Lub95, Ano96b]. Typography [Hos96a, Ric94, Ano96c, Lan94b, Sch96, Hos94a]. typophiles [Lan94a].

University [BM98, Ano96b].

vision [Lan95].

Walter [Ano95e]. WebType [BM98]. weeks [Bar95]. weight [Car94]. Wessex [Hos96d]. Westborough [BM98]. Which [Hos95i]. Whirligig [Ano94b]. Whither [Due95]. William [Pet94]. Windows [Ano94b]. wither [Due95]. Work [Hun98]. Works [Ric94].

x [Ano96b]. x-height [Ano96b].

Zimmerman [Pri95]. Zombie [Ano94a].

References

Anonymous:1994:BCF

[Ano94a] Anonymous. Big Caslon – FF Chelsea – Elroy – Lucida Casual – Twist – Zombie. Serif: The Magazine of Type and Typography, 1:6–9, Fall 1994. ISSN 1080-2207.

Anonymous:1994:BNF

[Ano94b] Anonymous. Briefly noted: Font CDs – Quickdraw GX – Multiple Masters for Windows – Ed Cleary – Treacyfaces and Headliners International – American Typecasting Fellowship – Emigre’s Whirligig Catalog – Santa Cruz Book Show. Serif: The Magazine of Type and Typography, 1: 10–12, Fall 1994. ISSN 1080-2207.

Anonymous:1995:BRF

[Ano95a] Anonymous. Book reviews: Fleischmann on Punchcutting. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207.

Anonymous:1995:BRL

[Ano95b] Anonymous. Book reviews: Letters of Credit. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207.

Anonymous:1995:BRT

[Ano95c] Anonymous. Book reviews: Twentieth Century Type Designers. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207.

Anonymous:1995:BNR

[Ano95d] Anonymous. Briefly noted: The return of ATF – ATypI ’94 – Recent publications – Type on the information superhighway – The death of the floppy? – News from the Goudy International Center. Serif: The Magazine of Type and Typography, 2:10–13, Winter 1995. ISSN 1080-2207.

Anonymous:1995:BNW

[Ano95e] Anonymous. Briefly noted: Walter Tracy – Toronto font party – Subiaco revival – Catalogs from Font Bureau and Precision Type – Bitstream lawsuit – Type on the information superhighway redux – Carl Herzog award – ATypI ’95 – FUSE ’95. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207.

Anonymous:1995:NFC

[Ano95f] Anonymous. New faces: Copal – FF Instanter – Magneto – Neonstream – Raceway – Rocket. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207.

Anonymous:1995:NFM

[Ano95g] Anonymous. New faces: Monotype Bulmer – Hairspray – Handsome – Schneidler Initials. Serif: The Magazine of Type and Typography, 2:7–9, Winter 1995. ISSN 1080-2207.

Anonymous:1995:T

[Ano95h] Anonymous. Technology. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207. Berthold’s Diamond CD.

Anonymous:1996:TDC

[Ano96a] Anonymous. 1996 typographic design competition: Information about entering the competition. Serif: The Magazine of Type and Typography, 4: 30–31, Spring 1996. ISSN 1080-2207.

Anonymous:1996:BNT

[Ano96b] Anonymous. Briefly noted: Typographic symposium and exposition — FontHaus, A Monotype company? — the new x-height — new Monotype CD — ITC expansion — Aldo Novarese – Monotype University — Typographic Circle debate — ATypI ’95 — the return of the bitmap. Serif: The Magazine of Type and Typography, 4:12–15, Spring 1996. ISSN 1080-2207.

Anonymous:1996:CCS

[Ano96c] Anonymous. A conversation with Carlos Segura: We discuss grunge typography, typographic collage, readability and originality with Carlos Segura. Serif: The Magazine of Type and Typography, 4:32–40, Spring 1996. ISSN 1080-2207.

Anonymous:1996:NFA

[Ano96d] Anonymous. New faces: Aries FH — Borek — Civilité MJ — Duckling — Jecho Techo One — Kalendar. Serif: The Magazine of Type and Typography, 4:8–11, Spring 1996. ISSN 1080-2207.

Anonymous:1998:NFB

[Ano98] Anonymous. New faces: Beata, Cresci, Donatello, Ghiberti, Pietra, Pontif, Simona. Serif: The Magazine of Type and Typography, 6:5–8, Spring 1998. ISSN 1080-2207.

Barrett:1995:CCG

[Bar95] Bonnie Barrett. Callifonts: Can a good font be made in two weeks? Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207. Barrett takes a hard look at the quality of the types in this broad and inexpensive collection.

Beaujon:1995:B

[Bea95] Beatrice Beaujon. In the beginning. Serif: The Magazine of Type and Typography, 2:46–51, Winter 1995. ISSN 1080-2207. A look at the use of initials in text showing the best way to work with both plain and decorated initials.

Beaujon:1996:DLP

[Bea96] Beatrice Beaujon. Designing for letter paper: Dealing with the design challenges of the 8.5 by 11 sheet of paper. Serif: The Magazine of Type and Typography, 4:46–51, Spring 1996. ISSN 1080-2207.

Bigelow:1994:TCS

[Big94] Charles Bigelow. Thoughts on Century Schoolbook. Serif: The Magazine of Type and Typography, 1:59– 60, Fall 1994. ISSN 1080-2207. A fresh look at a typeface we all have and seldom use.

Bruton:1998:BNA

[BM98] Clive Bruton and Rod McDonald. Briefly noted: ATypI Conference: 23–25 October 1998, Lyon France; TypeCon 98: 30 October– 1 November 1998, Westborough, Massachusetts; Alphabets, Inc. Restructures; Adobe WebType; Microsoft OpenType Font Jamboree; 1997 ATypI Conference, University of Reading. Serif: The Magazine of Type and Typography, 6:9–, Spring 1998. ISSN 1080-2207.

Bringhurst:1994:CL

[Bri94] Robert Bringhurst. On the classification of letterforms. Serif: The Magazine of Type and Typography, 1:30–39, Fall 1994. ISSN 1080-2207. An expansion and explication of the taxonomy begun in Bringhurst’s Elements of Typographic Style. First in a series.

Bringhurst:1995:HIH

[Bri95a] Robert Bringhurst. Holding ideas in the hand: The physics & metaphysics of renaissance letterforms. Serif: The Magazine of Type and Typography, 2: 15–24, Winter 1995. ISSN 1080-2207. A look at the first roman and italic typefaces, in the Montane, Humanist and Mannerist genera. Second in a continuing series.

Bringhurst:1995:PIS

[Bri95b] Robert Bringhurst. Painting with ink & steel: The roman & italic types of the baroque age. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207. An examination of the cultural context of the letters of the great baroque typographers including Jannon, Kis, Van Dijck and Caslon.

Bringhurst:1996:IHP

[Bri96] Robert Bringhurst. The invisible hand part one: Neoclassical letterforms: An examination of the development of the neoclassical letterforms of such designers as Baskerville, Bell and Fournier. Serif: The Magazine of Type and Typography, 4:19–29, Spring 1996. ISSN 1080-2207.

Carr:1994:WSO

[Car94] Dan Carr. The weight of a serif, or, the influences of punchcutting on type design. Serif: The Magazine of Type and Typography, 1:26–29, Fall 1994. ISSN 1080-2207. A look at the art of punchcutting and the degree to which it did—and didn’t—influence the development of letterforms.

Duensing:1995:MTW

[Due95] Paul Hayden Duensing. Metal type: Whither or wither? A look at the state of letterpress printing in 1994. Serif: The Magazine of Type and Typography, 2:25–27, Winter 1995. ISSN 1080-2207. A look at the current state of letterpress printing along with a list of current suppliers of metal type.

Frascara:1998:DEL

[Fra98] Jorge Frascara. Design education from lettering to human interaction. Serif: The Magazine of Type and Typography, 6:15–17, Spring 1998. ISSN 1080-2207.

Haley:1998:CCH

[Hal98] Allan Haley. Cool cash for hot fonts. Serif: The Magazine of Type and Typography, 6:19–22, 55, Spring 1998. ISSN 1080-2207.

Hosek:1994:BRM

[Hos94a] D. A. Hosek. Book reviews: Modern Typography, by Robin Kinross. Serif: The Magazine of Type and Typography, 1:64–65, Fall 1994. ISSN 1080-2207.

Hosek:1994:CA

[Hos94b] D. A. Hosek. Centaur & Arrighi. Serif: The Magazine of Type and Typography, 1:13–15, Fall 1994. ISSN 1080-2207. A look at the typeface used in the issue, its origins and development.

Hosek:1994:E

[Hos94c] D. A. Hosek. Endpapers. Serif: The Magazine of Type and Typography, 1: 68, Fall 1994. ISSN 1080-2207.

Hosek:1995:AP

[Hos95a] D. A. Hosek. Adobe poetica. Serif:

The Magazine of Type and Typography, 2:52–53, Winter 1995. ISSN 1080-2207. A review of Robert Slimbach’s Poetica.

Hosek:1995:CTSc

[Hos95b] D. A. Hosek. Contemporary type specimen books: Adobe Type Specimen Books. Serif: The Magazine of Type and Typography, 2:61–62, Fall 1995. ISSN 1080-2207.

Hosek:1995:CTSb

[Hos95c] D. A. Hosek. Contemporary type specimen books: American Metal Typefaces of the Twentieth Century, by Mac McGrew. Serif: The Magazine of Type and Typography, 2:59, 61, Fall 1995. ISSN 1080-2207.

Hosek:1995:CTSd

[Hos95d] D. A. Hosek. Contemporary type specimen books: An Atlas of Typeforms, by James Sutton and Alan Bartram. Serif: The Magazine of Type and Typography, 2:62–63, Fall 1995. ISSN 1080-2207.

Hosek:1995:CTSa

[Hos95e] D. A. Hosek. Contemporary type specimen books: M&H Type Catalogue. Serif: The Magazine of Type and Typography, 2:58–59, Winter 1995. ISSN 1080-2207.

Hosek:1995:CCT

[Hos95f] D. A. Hosek. A conversation with Carol Twombly. Serif: The Magazine of Type and Typography, 2:32– 40, Winter 1995. ISSN 1080-2207. A conversation with a type designer whose designs are seen everywhere and who was recently awarded the Prix Charles Peignot.

Hosek:1995:E

[Hos95g] D. A. Hosek. Endpapers. Serif: The Magazine of Type and Typography, 2: 64, Fall 1995. ISSN 1080-2207.

Hosek:1995:FLA

[Hos95h] D. A. Hosek. O: The fifteenth letter of the alphabet. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207.

Hosek:1995:WC

[Hos95i] D. A. Hosek. Which Caslon? Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207. A comparison of eight different digital Caslons and their suitability to different tasks.

Anonymous:1996:BRF

[Hos96a] D. A. Hosek. Book reviews: Finer Points in the Spacing & Arrangement of Type – The Form of the Book: Essays on the Morality of Good Design – The Art of Typography: An Introduction to Typo.icon.ography. Serif: The Magazine of Type and Typography, 4:60–62, Spring 1996. ISSN 1080-2207.

Hosek:1996:E

[Hos96b] D. A. Hosek. Endpapers. Serif: The Magazine of Type and Typography, 4: 64, Spring 1996. ISSN 1080-2207.

Hosek:1996:TIT

[Hos96c] D. A. Hosek. Technology: International typefounders CD. Serif: The Magazine of Type and Typography, 4: 58–59, Spring 1996. ISSN 1080-2207.

Hosek:1996:W

[Hos96d] D. A. Hosek. Wessex: A look at a book face designed by Matthew Butterick. Serif: The Magazine of Type and Typography, 4:16–17, Spring 1996. ISSN 1080-2207.

Hosek:1998:E

[Hos98] D. A. Hosek. Endpapers. Serif: The Magazine of Type and Typography, 6: 56, Spring 1998. ISSN 1080-2207.

Hunziker:1998:TWA

[Hun98] Hans-Jurg Hunziker. The typographic work of Adrian Frutiger. Serif: The Magazine of Type and Typography, 6:32–43, Spring 1998. ISSN 1080-2207.

Johnston:1995:TCT

[Joh95] Alastair Johnston. Two centuries of type & technology. Serif: The Magazine of Type and Typography, 2:28–31, Winter 1995. ISSN 1080-2207. A look back at bad taste in typography—it’s nothing new.

Keedy:1998:TC

[Kee98] Jeffery Keedy. A type of change.

Serif: The Magazine of Type and Typography, 6:44–48, 50–51, 55, Spring 1998. ISSN 1080-2207.

Lange:1994:DTF

[Lan94a] Gerald Lange. Digital type foundry for typophiles: Carter & Cone Type Inc. Serif: The Magazine of Type and Typography, 1:50–56, 58, Fall 1994. ISSN 1080-2207. A font-by-font accounting of the new designs by Matthew Carter as part of his new enterprise.

Lange:1994:TT

[Lan94b] Gerald Lange. [T-26] and ‘design typography’. Serif: The Magazine of Type and Typography, 1:46–48, Fall 1994. ISSN 1080-2207. A look at the output of one of a new breed of type foundries pushing the envelope of type design.

Lange:1995:GDG

[Lan95] Gerald Lange. Galápagos Design Group: Too much fashion—not enough vision. Serif: The Magazine of Type and Typography, 2:54–57, Winter 1995. ISSN 1080-2207. A look at the output of the newest Bitstream spin-off.

Long:1995:DKL

[Lon95] Robert Long. David Kindersley: In love with letters. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2207. A look at an artist whose lettering work has had perhaps less influence on typography than it should have.

Lubell:1995:TSF

[Lub95] Stephen Lubell. On typographic style: A few thoughts on legibility, readability and typographic tradition. Serif: The Magazine of Type and Typography, 3, Summer 1995.

ISSN 1080-2207. What makes for good typographic design? Lubell asks the questions and comes up with some answers.

Pastor:1994:BRA

[Pas94] Jon Pastor. Book reviews: Anatomy of a Typeface, by Alexander Lawson. Serif: The Magazine of Type and Typography, 1:65–66, Fall 1994. ISSN 1080-2207.

Peterson:1994:WMH

[Pet94] William S. Peterson. William Morris and his types. Serif: The Magazine of Type and Typography, 1:16–25, Fall 1994. ISSN 1080-2207. The original work and digital revivals as well.

Porchez:1998:H

[Por98] Jean-François Porchez. Horizontalité.

Serif: The Magazine of Type and Typography, 6:24–31, Spring 1998. ISSN 1080-2207.

Priven:1995:TTA

[Pri95] Aaron Priven. The times they are achanging (apologies to Mr. Zimmerman). Serif: The Magazine of Type and Typography, 2:45, Winter 1995. ISSN 1080-2207. A song about the new generation of type designers.

Rehak:1996:HGL

[Reh96] Theo Rehak. A hitchhiker’s guide to letterpress in the twenty-first century: An overview of methods for creating type for letterpress printing. Serif: The Magazine of Type and Typography, 4:42–45, Spring 1996. ISSN 1080-

Ricker:1994:BRH

[Ric94] William Ricker. Book reviews: How Typography Works, by Fernand Baudin. Serif: The Magazine of Type and Typography, 1:66, Fall 1994. ISSN 1080-2207.

Saxe:1994:ATS

[Sax94] Stephen O. Saxe. American type specimen books. Serif: The Magazine of Type and Typography, 1:61–63, Fall 1994. ISSN 1080-2207. A look at the development of a now seldom-seen product from type companies: The type specimen book.

Saxe:1995:MAT

[Sax95] Stephen O. Saxe. More about type specimens. Serif: The Magazine of Type and Typography, 2:42–44, Winter 1995. ISSN 1080-2207. A look at three noteworthy type specimen books.

Schwartz:1994:CG

[Sch94] Christian Schwartz. The changing of the guard. Serif: The Magazine of Type and Typography, 1:49, Fall 1994. ISSN 1080-2207. One type designer’s view of the changes in the business of type.

Schwartz:1996:HIT

[Sch96] Christian Schwartz. House industries: There goes the neighborhood: A look at a foundry producing grunge typography with a style of its own. Serif: The Magazine of Type and Typography, 4:53–57, Spring 1996. ISSN 1080-2207.

Solyn:1994:FIH

[Sol94] Paul Solyn. A fine italic hand. Serif:

The Magazine of Type and Typography, 1:40–45, Fall 1994. ISSN 1080-2207. A look at the history and use of italics.

Swanson:1994:CDC

[Swa94] Gunnar Swanson. Clarety: Drinking from the crystal goblet. Serif: The Magazine of Type and Typography, 1: 45, Fall 1994. ISSN 1080-2207. An irreverent look at whether Beatrice Warde’s metaphor still applies today.

VanBlokland:1998:ST

[Van98] Erik Van Blokland. The supermarket theory. Serif: The Magazine of Type and Typography, 6:52–55, Spring 1998. ISSN 1080-2207.

Wilmut:1998:LRC

[WR98] Sean Wilmut and Lindsay Rollo. Letters: Reader comments; on readability. Serif: The Magazine of Type and Typography, 6:4, Spring 1998. ISSN 1080-2207.